How do you solve this by using the addition or subtraction method? 8a+5b=9 2a–5b=–4
Answer by ankor@dixie-net.com(22740)

8a + 5b = 9
2a – 5b =–4
--------------Addition eliminates b, find a
10a = 5
a = 5/10
a = .5
:
Find b using the 1st equation
8(.5) + 5b = 9
4 + 5b = 9
5b = 9 - 4
5b = 5
b = 5/5
b = 1
:
Check solutions using the 2nd equation
2(.5) - 5(1) = -4
1 - 5 = -4